OpenStudy (anonymous):

If a line segment has the endpoints (2, 5) and (1, 1), what is its slope?

OpenStudy (nincompoop):

post question and run?

OpenStudy (anonymous):

\[ m = \frac{y_2-y_1}{x_2-x_1} = \frac{1-5}{1-2} \]

OpenStudy (anonymous):

Here I'm saying \[ (x_1,y_1)=(2,5)\quad(x_2,y_2)=(1,1) \]

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

medal